OpenStudy (anonymous):

factor the problem below x^2-10x+25

OpenStudy (anonymous):

\[(x-a)^2=x^2-2ax+a^2\] \[x^2-10x+25=(x-5)^2\]

OpenStudy (anonymous):

can you do x^2-3x-40

OpenStudy (anonymous):

yes.(x+5)(x-8)

OpenStudy (anonymous):

x^2-2x-15?

OpenStudy (anonymous):

(x-5)(x+3)